Oxygen ( molecular weight 32) is compressed reversibly and polytropically in a cylinder


from 1.05 bar, 15oC to 4.2bar in such a way that one- third of the work input is rejected


as heat to the cylinder walls . calculate the final temperature of the oxygen. Assume


oxygen to be a perfect gas and take Cv = 0.649kJ/kg.K

1
Expert's answer
2022-04-24T17:19:10-0400

"m=\frac{Mp_1V}{RT_1},"

"A=(p_2-p_2)V,"

"\frac 13 A=cm∆T,"

"T_2=T_1+\frac{(p_2-p_1)RT_1}{3cMp_1}=403~K=130°C."